OpenStudy (anonymous):

3 − 5(x + 2) = 3(x + 1) + 6 In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (aripotta):

First, distribute the -5 into the parentheses. 3 - 5x - 10 = 3(x + 1) + 6 Distribute the 3 into the parentheses. 3 - 5x - 10 = 3x + 3 + 6 Combine like terms. -5x - 7 = 3x + 9 Subtract 3x from both sides. -8x - 7 = 9 Add 7 to both sides. -8x = 16 Divide both sides by -8. x = -2